OpenStudy (anonymous):

Linda has two porcelain spheres. The radii of the spheres are in a ratio of 2 : 3. What is the ratio between the volumes of the two spheres?

OpenStudy (saifoo.khan):

\[\Large (\frac{L_1}{L_2})^3 = \frac{V_1}{V_2}\]

OpenStudy (anonymous):

Volume is 4/3*pi*r^3 so the volumes vary according to the cube of the radius